# ContextualStory: Consistent Visual Storytelling with Spatially-Enhanced and Storyline Context

### v2

@@ Introduction
Solving the Story Continuation Task
For story continuation tasks, besides the storyline, the first frame is provided as an extra input. As shown in Figure 4, we extract the first frame latent $\mathbf{z^{1}}$ and use it as an additional input to all UNet blocks. Within each UNet block, $\mathbf{z^{1}}$ is resized to match the spatial dimensions of the hidden state, then a $1\times 1$ convolution layer adjusts the channels to match the hidden state, and finally, it is concatenated with the hidden state before inputting into the spatial convolution.
Experiment
Experimental Setup
Datasets. We employ two popular benchmark datasets, PororoSV (Li et al. 2019) and the FlintstonesSV (Gupta et al. 2018), to evaluate the performance of our model in both story visualization and story continuation tasks. PororoSV contains 10,191, 2,334, and 2,208 stories within the train, validation, and test splits, respectively, including 9 main characters. On the other hand, FlintstonesSV contains 20,132, 2,071, and 2,309 stories within the train, validation, and test splits, respectively, featuring 7 main characters and 323 backgrounds. Each story within these datasets comprises 5 consecutive story images.
Automatic metrics. To evaluate the quality of generated images, we employ the following three evaluation metrics following previous works (Maharana, Hannan, and Bansal 2022; Pan et al. 2024) in story visualization: (1) Frechet Inception Distance (FID) (Heusel et al. 2017): Measures the distance between feature vectors of ground truth and generated frames. (2) Frame accuracy (Frm. Acc.): Evaluates character matching to ground truth using a fine-tuned Inception-v3 model. (3) Character F1-score (Char. F1): Evaluates the quality of generated characters using the same model as Frm. Acc.
Implementation details. We initialize ContextualStory with the pre-trained Stable Diffusion 2.1-base, fine-tuning only the UNet parameters using the AdamW optimizer. We use a batch size of 12, a learning rate of 5e-5, and train for 40,000 iterations on PororoSV and 80,000 iterations on FlintstonesSV, utilizing 4 NVIDIA A800 GPUs. The window size for SETA is set to $k=3$, and the number of SC layers is set to 4. During training, we apply classifier-free guidance by randomly dropping input text prompts with a 0.1 probability and use the PYoCo mixed noise prior for noise initialization. For inference, we use the DDIM sampler with 50 steps and a guidance scale of 7.5 to generate $256\times 256$ images.
Quantitative Results
Story Visualization. Table 2 shows quantitative results for story visualization on the PororoSV, comparing our ContextualStory to several SOTA methods, including StoryGAN, CP-CSV (Song et al. 2020), DUCO (Maharana, Hannan, and Bansal 2021), VLC (Maharana and Bansal 2021), VP-CSV (Chen et al. 2022), Word-Level SV (Li 2022), Story-LDM, AR-LDM, Causal-Story (Song et al. 2024) and StoryImager. Similarly, Table 4 shows quantitative results on the FlintstonesSV dataset. The quantitative results in Table 2 and Table 4 demonstrate that our ContextualStory significantly outperforms existing SOTA methods across all metrics on both datasets. This superior performance is primarily due to SETA, SC and StoryFlow Adapter, which effectively utilize context information to generate coherent story frames aligned with a given storyline.
Story Continuation. Table 2 presents the quantitative results for story continuation on both the PororoSV and FlintstonesSV datasets. We evaluate the effectiveness of our ContextualStory model against several SOTA methods, such as StoryDALL-E, MEGA-StoryDALL-E (Maharana, Hannan, and Bansal 2022), StoryImager, and RCDMs. The quantitative results in Table 2 demonstrate that ContextualStory outperforms existing methods by a large margin in all metrics for the story continuation task on both datasets. This indicates that our model can better utilize contextual information to generate a coherent story frames based on the given storyline and the first image.
Qualitative Results
Story Visualization. Figure 5 shows a qualitative comparison of story visualization on the PororoSV and FlintstonesSV. Stable Diffusion (SD) generates high-quality images independently based on individual sentences, but its lack of contextual awareness leads to inconsistent character appearances and character duplication. AR-LDM avoids character duplication but still struggles with inconsistent character appearances. In contrast, ContextualStory produces high-quality images with coherent and consistent characters and scenes across both datasets.
Story Continuation. Figure 11 demonstrates a qualitative comparison of story continuation on the PororoSV and FlintstonesSV datasets. StoryDALL-E produces low-quality characters with inconsistent backgrounds. AR-LDM generates higher-quality characters, but the backgrounds lack consistency and deviate significantly from the ground truth. In contrast, ContextualStory generates high-quality images with consistent characters and backgrounds that closely match the ground truth. More results are provided in the supplementary material.
Human Evaluation
Due to the limitations of metrics such as FID, Char. F1, and Frm. Acc. in accurately reflecting the quality of generated story frames, we conducted human evaluations for the story visualization task on the PororoSV and FlintstonesSV, focusing on Visual Quality, Semantic Relevance, and Temporal Consistency. We randomly selected 300 pairs of story frame sequences generated by AR-LDM (Pan et al. 2024) and our ContextualStory. Annotators were tasked to select better sequence for three attributes: Visual Quality, Semantic Relevance, and Temporal Consistency. Each pair of generated story frame sequences was evaluated by 10 annotators. As shown in Table 5, the evaluation results indicate that our ContextualStory outperforms AR-LDM significantly across all three attributes.
Figure 7: Qualitative results of ablation study on the proposed components for the story visualization task on the PororoSV.
Ablation Studies
Ablation study of the proposed components. To evaluate the benefit of each proposed component, we conduct an ablation study results of the proposed components for the story visualization task on the PororoSV. As shown in Table 4, progressively removing components from ContextualStory results in a decline across all three metrics. The removal of SETA has the most significant effect, increasing FID by 16.0%, and reducing Char. F1 and Frm. Acc. by 4.9% and 7.4%, respectively. The qualitative comparison in Figure 7 shows the following: 1) Removing the StoryFlow Adapter slightly reduces background consistency. 2) Further removing SC leads to duplicated characters, like Loopy. 3) Removing SETA reduces background consistency, introduces duplicated characters (e.g., Pororo), and incorrect characters (e.g., Petty and Poby), making images less accurate. 4) Removing Temporal Convolution further decreases character and scene consistency. These results indicate that all proposed components contribute to the performance of ContextualStory, with SETA having the most significant impact.
Ablation study of temporal attention. Table 6 presents the ablation study results of comparing Vanilla Temporal Attention and our proposed Spatially-Enhanced Temporal Attention for the story visualization task on the PororoSV. The results clearly show that SETA outperforms Vanilla Temporal Attention across all metrics. Specifically, SETA achieves a lower FID score, indicating better alignment with ground truth images, and higher Char. F1 and Frm. Acc., demonstrating improved character consistency and accuracy. These improvements highlight the effectiveness of the local window mechanism of SETA, which allows the model to better capture both spatial and temporal dependencies, leading to more coherent and consistent story frames. The significant gains in performance suggest that incorporating spatial context within temporal attention is crucial for enhancing visual storytelling models. More ablation studies are provided in the supplementary material.
@@
Method FID $\downarrow$ Char. F1 $\uparrow$ Frm. Acc. $\uparrow$
@@
Table 6: Ablation study of temporal attention for the story visualization task on the PororoSV.
@@ Conclusion
In this paper, we propose ContextualStory, a novel framework that overcomes the limitations of traditional autoregressive methods in visual storytelling. By incorporating Spatially-Enhanced Temporal Attention, we effectively capture spatial and temporal dependencies, ensuring consistency in characters and scenes across frames. Additionally, the Storyline Contextualizer enriches the global context from the entire storyline, while the StoryFlow Adapter enhances the model’s ability to handle scene changes. Extensive experiments on PororoSV and FlintstonesSV show that ContextualStory significantly outperforms state-of-the-art methods in both story visualization and continuation.
@@
Appendix A More Related Works
Method FID $\downarrow$ Char. F1 $\uparrow$ Frm. Acc. $\uparrow$
Simplified Storyline Contextualizer 13.94 74.60 47.18
Dual Self-Attention Storyline Contextualizer 13.75 75.31 47.84
Causal Storyline Contextualizer 14.59 75.02 48.00
Storyline Contextualizer (Ours) 13.61 77.24 51.59
Number of Layers FID $\downarrow$ Char. F1 $\uparrow$ Frm. Acc. $\uparrow$
1 13.96 75.20 48.26
2 13.84 76.06 49.56
4 (Ours) 13.61 77.24 51.59
8 14.78 74.73 46.88
Random Initialization 13.84 76.06 49.56
Zero Initialization (Ours) 13.61 77.24 51.59
Random Initialization 14.66 75.40 48.03
Zero Initialization (Ours) 13.61 77.24 51.59
Conv. First Image (Ours) 13.86 76.25 50.72
Concat. First Image 14.17 74.82 47.81
Appendix B UNet Architecture
@@
Appendix C Details of Training
@@
Appendix D Details of Datasets
@@
Appendix E More Ablation Studies
@@
Appendix F More Quantitative Results
Model Inference Speed
AR-LDM 40.4 s
StoryGen 31.7 s
ContextualStory (Ours) 11.8 s
Table 14: Inference speed comparison of different SOTA models.
Inference Speed. We compare the inference speed of different models, including AR-LDM (Pan et al. 2024), StoryGen (Liu et al. 2024), and our proposed ContextualStory, as shown in Table 14. The experiment is conducted on the same A800 GPU to ensure a fair comparison, with all models using 50 steps for the DDIM inference process. Both AR-LDM and StoryGen are autoregressive methods, which inherently suffer from slower inference speeds due to their sequential generation process. In contrast, ContextualStory is a non-autoregressive model, effectively addressing the inference speed bottleneck of autoregressive methods. The results demonstrate that ContextualStory achieves a significantly faster inference speed of 11.8 seconds, compared to 40.4 seconds for AR-LDM and 31.7 seconds for StoryGen. This indicates that ContextualStory not only offers a substantial improvement in inference speed over existing state-of-the-art methods but also outperforms them in terms of overall performance.
Appendix G More Qualitative Results
Appendix H Limitations

### v3

@@ Introduction
Model Memory (GB)$\downarrow$ Inference Speed (s)$\downarrow$
StoryDALL-E 20 347
Story-LDM 11 18.5
AR-LDM 40 40.4
StoryGen 29 31.7
StoryGPT-V 25 14.1
RCDMs 22 30.4
ContextualStory 5 11.8
Table 3: Comparison of GPU memory usage and inference speed across SOTA models.
@@
Figure 7: Qualitative results of the ablation study on the proposed components for story visualization on PororoSV.
Quantitative Results
Story Visualization. Table 2 presents quantitative results for story visualization on both PororoSV and FlintstonesSV, comparing ContextualStory to several SOTA methods, including StoryGAN, CP-CSV (Song et al. 2020), DUCO (Maharana, Hannan, and Bansal 2021), VLC (Maharana and Bansal 2021), VP-CSV (Chen et al. 2022), Word-Level SV (Li 2022), Story-LDM, AR-LDM, Causal-Story (Song et al. 2024), and StoryImager. The results clearly demonstrate that ContextualStory significantly outperforms existing SOTA methods across all metrics on both datasets. This superior performance is primarily due to SETA, SC, and StoryFlow Adapter, which effectively utilize context information to generate coherent story frames.
Story Continuation. Table 1 presents the quantitative results for story continuation on both PororoSV and FlintstonesSV. We evaluate the effectiveness of ContextualStory model against several SOTA methods, such as StoryDALL-E, MEGA-StoryDALL-E (Maharana, Hannan, and Bansal 2022), StoryImager, and RCDMs. The results demonstrate that ContextualStory outperforms existing methods by a large margin across all metrics for the story continuation on both datasets. This indicates that ContextualStory better utilize contextual information to generate coherent story frames based on the storyline and the first image.
Inference Speed. We compare the GPU memory usage and inference speed of recent open-source SOTA models (e.g., AR-LDM and StoryGen) in Table 3. The experiment is conducted on an A800 GPU with 50 DDIM steps to ensure a fair comparison. Autoregressive methods like Story-LDM, AR-LDM, StoryGen, and StoryGPT-V suffer from high memory usage and slow inference speeds. In contrast, ContextualStory, a non-autoregressive model, not only overcomes the bottleneck of autoregressive methods by achieving the lowest memory usage and fastest inference speed, but also outperforms SOTA methods in overall performance.
Qualitative Results
Story Visualization. Figure 5 shows a qualitative comparison of story visualization on PororoSV and FlintstonesSV. Stable Diffusion (SD) generates high-quality images independently from individual sentences, but its lack of contextual awareness leads to inconsistent character appearances and character duplication. AR-LDM avoids character duplication but still struggles with inconsistent character appearances. In contrast, ContextualStory produces high-quality images with coherent and consistent characters and scenes across both datasets.
Story Continuation. Figure 10 demonstrates a qualitative comparison of story continuation on PororoSV and FlintstonesSV datasets. StoryDALL-E produces low-quality characters with inconsistent backgrounds. AR-LDM generates higher-quality characters, but the backgrounds lack consistency and deviate significantly from the ground truth. In contrast, ContextualStory generates high-quality images with consistent characters and backgrounds that closely match the ground truth. More results are provided in the supplementary material.
Human Evaluation
Due to the limitations of metrics such as FID, Char. F1, and Frm. Acc. in accurately reflecting the quality of generated story frames, we conducted human evaluations for the story visualization task on PororoSV and FlintstonesSV, focusing on Visual Quality, Semantic Relevance, and Temporal Consistency. We randomly selected 300 pairs of story frame sequences generated from AR-LDM (Pan et al. 2024) and our ContextualStory. Annotators were tasked to select the better sequence for the three attributes: Visual Quality, Semantic Relevance, and Temporal Consistency. Each pair of story frame sequences was evaluated by 10 annotators. As shown in Table 4, the results indicate that our ContextualStory outperforms AR-LDM significantly across all three attributes.
Ablation Studies
Ablation study of the proposed components. To evaluate the benefit of each proposed component, we conduct an ablation study on the story visualization task using PororoSV. As shown in Table 5, progressively removing components from ContextualStory results in a consistent decline across all three metrics. The removal of SETA has the most significant effect, increasing FID by 16.0%, and reducing Char. F1 and Frm. Acc. by 4.9% and 7.4%, respectively. The qualitative comparison in Figure 7 shows the following: 1) Removing the StoryFlow Adapter slightly reduces background consistency. 2) Further removing SC leads to duplicated characters, like Loopy. 3) Removing SETA reduces background consistency, introduces duplicated characters (e.g., Pororo), and incorrect characters (e.g., Petty and Poby), making images less accurate. 4) Removing Temporal Convolution further decreases character and scene consistency. These results indicate that all proposed components contribute to the performance of ContextualStory, with SETA having the most significant impact.
Ablation study of temporal attention. Table 5 presents the ablation study results of comparing Vanilla Temporal Attention and our proposed SETA for the story visualization task on PororoSV. The results clearly show that SETA outperforms Vanilla Temporal Attention across all metrics. Specifically, SETA achieves a lower FID score, indicating better alignment with ground truth images, and higher Char. F1 and Frm. Acc., demonstrating improved character consistency and accuracy. These improvements highlight the effectiveness of the local window mechanism of SETA, which allows the model to better capture both spatial and temporal dependencies, leading to more coherent and consistent story frames. The significant gains in performance suggest that incorporating spatial context within temporal attention is crucial for enhancing visual storytelling models. More ablation studies are provided in the supplementary material.
Model FID $\downarrow$ Char. F1 $\uparrow$ Frm. Acc. $\uparrow$
ContextualStory 13.61 77.24 51.59
--StoryFlow Adapter 14.84 77.09 50.48
     --Storyline Contextualizer 15.02 75.42 48.39
      --SETA 17.42 71.70 44.83
       --Temporal Convolution 19.69 68.12 39.60
@@
Table 5: Ablation study of the proposed components and temporal attention for story visualization on PororoSV.
@@ Conclusion
In this paper, we propose ContextualStory, a novel framework that overcomes the limitations of traditional autoregressive methods in visual storytelling. By incorporating Spatially-Enhanced Temporal Attention, we effectively capture spatial and temporal dependencies, ensuring consistency in characters and scenes across frames. Additionally, the Storyline Contextualizer enriches the global context from storyline, while the StoryFlow Adapter enhances the ability of model to handle scene changes. Extensive experiments on PororoSV and FlintstonesSV show that ContextualStory achieves superior performance over state-of-the-art methods in both story visualization and continuation tasks.
Appendix A Acknowledgments
This paper is supported by the Doubao Fund.
@@ Appendix B More Related Works
Appendix C UNet Architecture
@@
Appendix D Details of Training
@@
Appendix E Details of Datasets
@@
Appendix F More Quantitative Results
Method FID $\downarrow$ Char. F1 $\uparrow$ Frm. Acc. $\uparrow$
Simplified Storyline Contextualizer 13.94 74.60 47.18
Dual Self-Attention Storyline Contextualizer 13.75 75.31 47.84
Causal Storyline Contextualizer 14.59 75.02 48.00
Storyline Contextualizer (Ours) 13.61 77.24 51.59
Appendix G More Qualitative Results
Appendix H More Ablation Studies
@@
Number of Layers FID $\downarrow$ Char. F1 $\uparrow$ Frm. Acc. $\uparrow$
1 13.96 75.20 48.26
2 13.84 76.06 49.56
4 (Ours) 13.61 77.24 51.59
8 14.78 74.73 46.88
Random Initialization 13.84 76.06 49.56
Zero Initialization (Ours) 13.61 77.24 51.59
Random Initialization 14.66 75.40 48.03
Zero Initialization (Ours) 13.61 77.24 51.59
Conv. First Image (Ours) 13.86 76.25 50.72
Concat. First Image 14.17 74.82 47.81
@@
Appendix I Limitations